The following data pertain to products A and B, both of which are purchased by Madame X. Initially, the prices of the products and quantities consumed are:

PA = $10, QA = 3, PB = $10, QB = 7.

Madame X has $100 to spend per time period. After a reduction in price of B, the prices and quantities consumed are:

PA = $10, QA = 2.5, PB = $5, QB = 15.

Assume that Madame X maximizes utility under both price conditions above. Also, note that if after the price reduction enough income were taken away from Madame X to put her back on the original indifference curve, she would consume this combination of A and B:

QA = 1.5, QB = 9

a. Determine the change in consumption rate of good B due to (1) the substitution effect and (2) the income effect. b. Determine if product B is a normal, inferior, or Giffen good. Explain.
Answer #1

A) total change in Consumption of Good B= 15-7 = 8

So total effect = 8

Now, substitution effect is found , utility level is kept constant, & we move along the original IC only

So SE = 9-7 = 2

As TE = SE + IE

so IE = 8-2 = 6

b) now as price of good B falls, then consumption of Good B rises , hence product B is normal (ordinary) good.